OpenStudy (anonymous):

any one who can help me with calc 3

OpenStudy (lgbasallote):

is calc 3 differential equations?

OpenStudy (anonymous):

F=cos(pi y i)-pi x sin(pi y)j ,,is this a gradient?

OpenStudy (anonymous):

Hmm it's a bit hard to read, for me. \[ F= \cos(\pi y)i -(\pi x \sin (\pi y)j \] ? I hardly doubt that I got it correct, so maybe you can draw it or upload a picture/copy of the problem?

OpenStudy (anonymous):

you got it right

OpenStudy (anonymous):

Hehe, the power of \( \LaTeX\), okay let me take a look at it.

OpenStudy (anonymous):

To verify if this is an exact equation you want to check if \[\Large M_y=N_x\] You might have seen that before in exact equations.

OpenStudy (anonymous):

Excuse me if I did empathize a bit too much with the exact equation terminology, it's how I recall it. Basically you use the same methods of checking, if the statements hold, you have a gradient field.

OpenStudy (anonymous):

you right,,,but i want to verify for that particular question,,

OpenStudy (anonymous):

is it gradient or not,,i got it's not but the answer turns to be it's gradient in my book

OpenStudy (anonymous):

Good, so maybe the following will do the trick. \[ \Large F= \underbrace{\cos(\pi y)}_Mi -\underbrace{(\pi x \sin (\pi y)}_Nj \]

OpenStudy (anonymous):

yes it is a gradient field, differentiate the M term with respect to y, partial derivative, then take the partial with respect to x of the second term.

OpenStudy (anonymous):

cos pi y gives me - pi( sin pi Y)

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

exactly \[\Large M_y=- \pi \sin (\pi y) \]

OpenStudy (anonymous):

the second piece didn't gave me the same answer though,,,any clue?

OpenStudy (anonymous):

\[\Large N= -\pi x \sin (\pi y) \] ( !! I did distribute the minus sign infront of the \(j\) component already))

OpenStudy (anonymous):

Now you want to take the partial of that with respect to x, so all the terms that include a y are just a constant.

OpenStudy (anonymous):

in fact \( \sin (\pi y) \) is a constant.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

really apperitiate

OpenStudy (anonymous):

very welcome, hope that helped.

OpenStudy (anonymous):

yep it did